OpenStudy (poopsiedoodle):

Does anyone here take bio with flvs that can help me on 4.03? I'm rather confused on that one.

OpenStudy (blues):

Poopsie, what is 4.03? It's a number, plainly, but the rest of us don't know what that corresponds to in terms of unit material...

OpenStudy (poopsiedoodle):

It's a lab on population dynamics.

OpenStudy (abb0t):

rate at which a population increases in size, (assuming there r no density-dependent forces regulating the population) is: intrinsic rate of increase \[\frac{ dN }{ dt }\frac{ 1 }{ N } = r\] There's also: \[x(t) = ab^{\frac{ t }{ \tau }} \] a is the initial value of x (meaning x(0) = a b is a positive growth factor, tau is the time constant (the time required for x to increase by one factor of b). There's also (if this is for a upper division biology course that involves differentials), you might use: \[\frac{ dx }{ dt } = kx\] x(0) = a a very simple and basic linear D.E: \[\int\limits \frac{ dx }{ x } = k dt\] ln(x) = kt + C and incorporating the initial values, x(0) = a thus, ln(x) = kt + ln(a) and finally, \[x = ae^{kt}\] and applies for a ≤ 0 It's different for a nonlinear DE, but I'll leave it at is for now. If you need more help, just ask.

OpenStudy (frostbite):

You can also do a logistic differential equation if working with a bacteria population with limit resources and different environment.

OpenStudy (poopsiedoodle):

I actually just need help on the part where it asks me for a hypothesis. I have no idea what I'm supposed to be hypothesizing about.

OpenStudy (anonymous):

UGH YES IM TAKING AP BIO ON FLVS IT SUCKKKKKKSSS

OpenStudy (abb0t):

oh

OpenStudy (anonymous):

I take bio in FLVS. if you ever need help I would be more than happy to. Just email me if you ever have any more questions. I help many others in Bio, and English.

OpenStudy (poopsiedoodle):

Well, if you could help me figure out what I'm supposed to hypothesize about, that'd be A+.

OpenStudy (anonymous):

Okay. Is this biology 1 or 2?

OpenStudy (poopsiedoodle):

Bio 1. I didn't know there was a Bio 2. I thought the 1 was just there because of reasons o-o

OpenStudy (anonymous):

No. I think that there is a bio 2. and i got at 100% on that assignment. And ican not exactly tell you what to hypothesis, but what i can do is give you what i said for the hypothesis, and you can either use that or put it into your own words.

OpenStudy (anonymous):

Note: Read the procedures section before completing your hypotheses. 1. Part One: 1. Seal population 2,238 seal carrying capacity 290 whale population 20,000 whale carrying capacity 1,600. 2. Seal Population 2,238 Seal Carrying Capacity 250 Whale Population 15,000 Whale Carrying Capacity 1,300. 2. Part Two: 1. Seal growth rate 3% whale death rate 42% seal carrying capacity 250 whale carrying capacity 1,500. 2. Seal growth rate 15% whale death rate 50% seal carrying capacity 285 whale carrying capacity 1,400. 3. Part Three: (this is what i put under hypothsis)

OpenStudy (poopsiedoodle):

Okay, because I have no idea what I'm doing :P

OpenStudy (anonymous):

Its all good. no worries. :)

OpenStudy (anonymous):

is that all hun?

OpenStudy (poopsiedoodle):

For now :P

OpenStudy (anonymous):

7. Climate change (an abiotic factor) has slowly decreased habitat for the Arctic seals. Predict how a reduction in habitat could change the capture efficiency of the killer whales in our simulation? How might this affect the population of both the whales and the seals?

OpenStudy (anonymous):

i need help with this question